l. For which of the following d will current flow through the light bulb? (all coils are complete closed circuit) 2. A bar magnet is falling through a loop of wie with constant vekeity.The south pole enters first. As the south pole of the magnet approaches the wire, the induced cument (as viewed from is clockwise. kwise is along the length of the magnet 3. The speed of ultraviolet radiation in vacuum compared with the speed of visible light in a. very much greater. d a little less. b, a little greater. e, very much less. the same 4. The wavelength of 200-MHz television signal is approvimately: b, 1,5 m 2.0m d, 2.0 cm
Answer #1

1 ans

the current passing 1 and 2

becasue the magnet is placed closed the loop the current is passing but in the loop 3 and 4 the current is does not passed because there is no change of flux linked with the coil

---------------------------------------------------------------------------------------------------------------------------------------------------------------

2 ans

basing on the faradays law the south pole entre to first in the loop the current is induced clockwise

the correct option B

-----------------------------------------------------------------------------------------------------------------------------------

3 ans

speed of UV light in vacum=2.9934*10^8 m/s

speed of visble light in vaccum =2.998*10^8 m/s

basing on this correct option is D

-----------------------------------------------------------------------------------------------------------------------------------------

4 ans

frequency f=200 Mhz=200*10^6 Hz

now we find the wavelength

wavelength=3*10^8/2*10^8=1.5 m

the correct option is B